OpenStudy (anonymous):

Solve the systems using substitution. x + 3y = -5 2x - y = 18

OpenStudy (anonymous):

3/ 2x-y=18, 6x -3y = 54, x+3y=-5 7x=49 x=7,

OpenStudy (anonymous):

x=7 y=-4